Roy's Welding common stock sells for $58.96 a share and pays an annual dividend that increases by 4.5 percent annually. The market rate of return on this stock is 16.5 percent. What is the amount of the dividend expected to be paid one year from today (next year)?

A.

$6.77

B.

$5.64

C.

$7.08

D.

$4.80
Answer #1

Stock price(P) = $58.96

Annual dividend = D

Growth rate(g) = 4.5%

Market rate of return (K) = 16.5%

P = [D(1+g) /(k-g)]

$58.96 = D(1.045) / (0.165 - 0.045)

$58.96 = D(1.045) / 0.12

D= 6.770526

Dividend expected to be paid one year from today is $6.770526 * (1+g)

D1 = 6.770526 * 1.045

D1 = $ 7.08

Option 'C' is correct